OpenStudy (anonymous):

What is the equation of the line that passes through (-2, 2) and (1, -4)? y = 2x - 2 y = -2x - 4 y = -2x - 2 y = 2x + 4

OpenStudy (anonymous):

is it y = -2x - 2

OpenStudy (shinalcantara):

slope formula: \[m = \frac{ y2-y1 }{ x2-x1 }\] \[m = \frac{ -4-2 }{ 1-(-2) }\] solve for the slope first

OpenStudy (anonymous):

-6/3= -2

OpenStudy (shinalcantara):

we'll use again that same formula with slope as -2 and a point. either (-2,2) or (1,-4) in this case let me use (1,-4) \[m = \frac{ y2-y1 }{ x2-x1 }\] \[-2 = \frac{ y - (-4) }{ x-1 }\] cross multiply \[-2(x-1) = y+4\] \[-2x + 2 = y + 4\] \[y + 2x + 4 - 2 = 0\] \[y + 2x + 2 = 0\] or in slope intercept form: \[y = -2x -2\]